OpenStudy (anonymous):

find an equation of line with slope -3/5 and goes through point (1,0) write equation using function notation

OpenStudy (yttrium):

Say we have m = -3/5 passing thru pt (1,0). Using point-slope form, we have. \[y-y _{1} = m (x-x _{1})\] Substituting the values, we'll have \[y-0 = \frac{ -3 }{ 5 } (x-1)\] Since we are required to write the equation using function notation, we'll let f(x) = y Hence: \[f(x) = \frac{ -3 }{ 5 } (x-1)\]

OpenStudy (anonymous):

so would it be that or would you go a step further and put f(x)=-3/5x+1

OpenStudy (yttrium):

That would be better indeed. :))

OpenStudy (anonymous):

ok thank you!